# Group-Adaptive Threshold Optimization for Robust AI-Generated Text Detection

### v4

@@ 4.4 Discovery of Subgroup Bias in Text Detection
We assess detection performance across text-length subgroups under a static threshold $\theta=0.5$. As illustrated in Figure 3, short human-authored samples incur a substantially higher FPR than mid-length and long texts. This length-dependent disparity underscores the necessity of text length adaptive thresholding to mitigate the discrepancy.
@@
Next, we examine the combined effects of length and stylistic personality (Figure 5). By comparing the empirical cumulative distribution functions (CDFs) of AI-likelihood scores for short, conscientious versus long, neurotic texts, a two-sample Kolmogorov–Smirnov (KS) test reveals a statistically significant divergence in their score distributions ($p<0.05$), indicating that static thresholding disproportionately misclassifies certain stylistic and length combinations (see Appendix G.1.1 for full KS-test details). The observed disparity in the combination of text length and stylistic personality further highlights the need for adaptive thresholding to guarantee equitable detection performance across heterogeneous text cohorts.
4.5 Mitigating Subgroup Bias using FairOPT
@@
Given these findings, we now describe the experimental setup for applying FairOPT for AI-text detection with reduced discrepancy across subgroups. We extracted the text length and personality features from every instance in the dataset, partitioning the corpus into 15 subgroups. For the dataset, multiple AI text classifiers were employed to generate predicted probabilities indicating the probability that a given text was AI-generated with $p\in[0,1]$.
@@
We used FairOPT to get adaptive thresholds $\{\theta(G_{1}),\dots,\theta(G_{j})\}$ for each subgroup and applied them as noted in Figure 2. We include all hyperparameters and training details in Appendix H.3. Threshold optimization was performed on the training dataset based on the results from RoBERTa-large [16] rather than averaging the probabilities of multiple detectors. The relaxed fairness criterion is integrated to encourage earlier convergence and minimize the sacrifice of overall performance while pursuing group fairness [46].
@@
We also applied an AUROC-based thresholding method using the RoBERTa-large [16] to select the decision threshold that maximizes the TPR under the constraint $FPR<0.1$; see Figure 6 for the AUROC-based optimization procedure [26]. The optimized thresholds were then applied to the test dataset and evaluated based on discrepancy, ACC, and F1 score. Metrics are noted in Appendix B.2.
4.6 Tradeoff Analysis between Performance and Fairness
@@
Figure 4 shows the performance-fairness tradeoff by our proposed FairOPT algorithm compared to the Static and ROCFPR baselines. The relaxed fairness criterion from Section 3.2 to less sacrificing ACC while still satisfying a demanded level of fairness.
@@
4.7 Results with FairOPT for AI-Text Detection
Figure 7 demonstrates that group-adaptive thresholding from FairOPT consistently lowers the average discrepancy. Our method decreased the discrepancy to $0.4341$ ($\sigma=0.1927$), showing $12\%$ of reduction compared to the static-threshold baseline ($\mu=0.5512$, $\sigma=0.1758$) and outperforms the AUROC-tuned method ($\mu=0.5595$, $\sigma=0.1582$). Notably, E5LoRA shows the greatest decrease, with its BER discrepancy dropping by nearly half (from $0.63$ to $0.33$), and GPT4o-mini achieved a substantial reduction (from $0.60$ to $0.37$). The Desklib experiences a modest 0.003 reduction. Overall, the FairOPT decreased the discrepancy in all AI text classifiers. The detailed results with standard deviation by the dataset, detector, and thresholding methods are in Appendix LABEL:tab:overall_by_detector_and_dataset.
Figure 8 illustrates that FairOPT incurs only a 0.003 of sacrifice in overall ACC and reduction of the standard deviation by 0.04 while outperforming the AUROC-based method. These findings demonstrate that FairOPT substantially mitigates fairness disparities with negligible ACC tradeoff, offering a practical application. Notably, our method shows great effectiveness to commonly used AI detectors like DeskLib [39], FastDetectGPT [41], RADAR [13], and SimpleAI’s detector [42] by securing ACC over $75\%$ while reducing the discrepancy in all cases.
Moreover, the E5LoRA detector [40] exemplifies the ability of FairOPT to transcend the fairness–performance tradeoff. On the SemEval24 dataset, FairOPT achieves a 0.43 reduction in discrepancy while increasing ACC by 0.23. Similarly, on the MAGE, it delivers a 0.25 decrease in discrepancy coupled with a 0.21 ACC increase. These outcomes highlight the broad applicability of adaptive thresholding for improving both fairness and performance across diverse AI-text detection models, as well as other kinds of probabilistic classifiers.
5 Conclusion
@@
Although our experiments focus on text-based detectors, the FairOPT framework is modality-agnostic and readily extends to other domains such as image, video, and audio deepfake detection, where probabilistic outputs require post-processing to ensure equitable performance. More generally, FairOPT applies to any probabilistic classification task with subgroup disparities (e.g., recidivism risk prediction, hiring decisions) by learning tailored decision thresholds that mitigate bias with minimal impact on overall accuracy.
6 Limitations
While FairOPT demonstrably reduces subgroup discrepancy on our English-text benchmarks, it relies on only two features (character-count bins and model-inferred Big Five traits). Other text attributes (e.g., topic, syntax, language) may induce additional biases that our current feature set does not capture. Moreover, the stylistic labels are derived from a downstream BERT-based personality predictor, which itself may introduce cultural bias. We focus on relaxing demographic parity—extensions to other fairness definitions (equal opportunity or equalized odds) may require different optimization objectives or could be fundamentally incompatible.
Impact Statement

### v5

@@ IV-D Discovery of Subgroup Bias in Text Detection
We assess detection performance across text‐length subgroups under a static threshold $\theta=0.5$. As illustrated in Appendix Detailed Results, short human‐authored samples incur a substantially higher FPR than mid‐length and long texts. This length-dependent disparity underscores the necessity of text length adaptive thresholding to mitigate the discrepancy.
@@
Next, we examine the combined effects of length and stylistic personality (Figure 5). By comparing the empirical cumulative distribution functions (CDFs) of AI‐likelihood scores for short, conscientious versus long, neurotic texts, a two‐sample Kolmogorov–Smirnov (KS) test reveals a statistically significant divergence in their score distributions ($p<0.05$), indicating that static thresholding disproportionately misclassifies certain stylistic and length combinations (see Appendix -F1 for full KS‐test details). The observed disparity in the combination of text length and stylistic personality further highlights the need for adaptive thresholding to guarantee equitable detection performance across heterogeneous text cohorts.
IV-E Mitigating Subgroup Bias using FairOPT
@@
Given these findings, we now describe the experimental setup for applying FairOPT for AI-text detection with reduced discrepancy across subgroups. We extracted the textual features from every instance in the dataset, partitioning the corpus into 60 subgroups. Across each dataset, multiple AI text classifiers were used to infer the probability $p\in[0,1]$ that a given text was AI-generated.
@@
We used FairOPT to get adaptive thresholds $\{\theta(G_{1}),\dots,\theta(G_{j})\}$ for each subgroup and applied them as noted in Figure 9. We include all hyperparameters and training details in Appendix -I. Threshold optimization was performed on the training dataset based on the results from RoBERTa-large [42] rather than averaging the probabilities of multiple detectors. The relaxed fairness criterion is integrated to encourage earlier convergence and minimize the sacrifice of overall performance while pursuing group fairness [44].
@@
We also applied an AUROC‐based thresholding method using the RoBERTa‐large [42] to select the decision threshold that maximizes the TPR under the constraint $FPR<0.1$; see Figure 6 for the AUROC‐based optimization procedure [36]. The optimized thresholds were then applied to the test dataset and evaluated based on discrepancy, ACC, and F1 score. Metrics are noted in Appendix -B.
IV-F Tradeoff Analysis between Performance and Fairness
@@
Figure 4 shows the performance-fairness tradeoff by our proposed FairOPT algorithm compared to the Static and ROCFPR baselines. The relaxed fairness criterion from Section III-B to less sacrificing ACC while still satisfying a demanded level of fairness.
@@
IV-G Results with FairOPT for AI-Text Detection
Table I shows that group-adaptive thresholding with FairOPT consistently reduces subgroup discrepancy across detectors. Overall, our method decreased the average discrepancy around $27.4\%$, substantially reducing subgroup gaps compared to static and ROCFPR baselines. On average, FairOPT achieves a BER discrepancy of $0.5852$, representing a $22\%$ reduction relative to the static threshold baseline ($0.7507$) and a $23\%$ reduction compared to ROCFPR ($0.7592$). Similar improvements are observed across other metrics: demographic parity decreases from $0.9730$ to $0.6772$ ($30\%$), equalized odds from $0.8501$ to $0.6379$ ($25\%$), false-positive rate discrepancy from $0.9471$ to $0.4313$ ($51\%$), and predictive parity from $0.9130$ to $0.7333$ ($18\%$). Overall, FairOPT lowers the discrepancy. Full results, including means and standard deviations by dataset, detector, and thresholding method, are provided in Appendix LABEL:tab:overall_by_detector_and_dataset.
Regarding the consistency of FairOPT across datasets, it successfully reduces discrepancy across all fairness metrics compared to both Static and ROCFPR, as shown in Table II. On average, the MaxBER discrepancy decreases by $22\%$, dropping from 0.751 (Static) and 0.759 (ROCFPR) to 0.585. Similarly, the MaxDP discrepancy decreases by $30\%$, from 0.973 (Static) and 0.983 (ROCFPR) to 0.677, while the MaxEO discrepancy decreases by $25\%$, from 0.850 (Static) and 0.920 (ROCFPR) to 0.638. The strongest improvement is observed for MaxFPR, where FairOPT achieves a $55\%$ reduction, falling from 0.947 (Static) and 0.972 (ROCFPR) to 0.431. Predictive Parity discrepancy also improves by $20\%$, decreasing from 0.913 (Static) and 0.968 (ROCFPR) to 0.733. Importantly, these fairness gains come with negligible accuracy cost: overall accuracy remains stable at 0.634 for FairOPT, essentially unchanged from 0.638 under Static and higher than 0.625 under ROCFPR.
Across detectors, FairOPT consistently reduces the average discrepancy as shown in 2 while making a negligible sacrifice of ACC 7. Across detectors, FairOPT lowers the mean discrepancy to 0.613 versus 0.887 (static) and 0.920 (ROCFPR). E5LoRA shows the largest gain, dropping to 0.100 from 0.672 (static) and 0.878 (ROCFPR). Among BERT-based classifiers, RoBERTa-base falls to 0.782 from 0.927 (static) and 0.932 (ROCFPR), while RoBERTa-large drops to 0.490 from 0.890 (static) and 0.940 (ROCFPR), both indicating substantial reductions. GPT-based detectors also improve meaningfully: GPT4o-mini to 0.253 and GPT4.1-nano to 0.466, compared to 0.9 under both static and AUROC. Other models show smaller but positive gains—DeskLib 12%, FastDetectGPT 10%, and SimpleAI 5%. The lone exception is RADAR, which increases by 2% under FairOPT; here, AUROC is approximately 1% lower than static. Detailed results with standard deviations are included in Appendix LABEL:tab:overall_by_detector_and_dataset.
Figure 7 shows that the fairness improvements achieved by FairOPT come with only negligible predictive cost. Averaged across all datasets, FairOPT achieves an accuracy of $0.6341$, compared to $0.6376$ for Static and $0.6249$ for ROCFPR. The difference compared to static is less than $0.1\%$. These findings confirm that FairOPT substantially mitigates fairness disparities with negligible accuracy tradeoff, making it a practical approach for fair AI-text detection.
IV-H Threshold Stability
The stability of FairOPT were systematically evaluated across three highly diverse and representative benchmark datasets: MAGE, RAID, and SemEval. These datasets were intentionally selected to capture a broad spectrum of domains, genres, and writing styles, including academic and encyclopedic texts, social media posts, creative writing, and reviews. This diversity ensures that our subgroup definitions, based on features such as length, personality, formality, and sentiment, are well supported and meaningful within each dataset. This provides a rigorous testbed for threshold optimization.
V Conclusion
@@
The novelty of this work resides within introducing a group-adaptive, post-hoc threshold optimization framework that operates without retraining of the detectors in the given settings by leveraging fairness and performance metrics. To best of our knowledge, no prior research has systematically examined threshold optimization in the context of AI text detectors although some research investigated thresholding methods for probabilistic classifiers. FairOPT represents a significant step toward AI safety by providing an alternative to the common practice of applying a single fixed threshold (e.g., 0.5, 0.75 or 0.97151).
@@
VI Limitations
While FairOPT demonstrably reduces subgroup discrepancy on our English‐text benchmarks, it relies on only two features (character‐count bins and model‐inferred Big Five traits). Other text attributes (e.g., topic, syntax, language) may induce additional biases that our current feature set does not capture. Moreover, the stylistic labels are derived from a downstream BERT‐based personality predictor, which itself may introduce cultural bias. We focus on relaxing demographic parity—extensions to other fairness definitions (equal opportunity or equalized odds) may require different optimization objectives or could be fundamentally incompatible.
VII Future Works
Although our experiments focused on English-language AI-generated text detectors, FairOPT is a modality-agnostic, subgroup-specific thresholding framework that readily extends to new deployment contexts and populations: for previously unseen subgroups, thresholds can be recalibrated from available data or warm-started from similar existing groups (e.g., AI-written text in schools, AI-generated images on a platform, or deepfakes on YouTube). Beyond text, FairOPT applies to image, video, and audio deepfake detection—any setting with probabilistic outputs that benefits from post-processing to ensure equitable performance—and more broadly to probabilistic classification tasks with subgroup disparities (e.g., recidivism risk prediction, hiring decisions), where learning tailored decision thresholds can mitigate bias with minimal impact on overall accuracy. We recognize that both theoretical analysis and user-centered case studies will be important for validating these extensions across diverse, real-world deployments.
VIII Impact Statement